OpenStudy (anonymous):

Exponential function help!! Can someone help me understand this? 8 = 2^(x+4)

OpenStudy (kohai):

What do you need to do with this function?

OpenStudy (anonymous):

I need to solve for x.

OpenStudy (kohai):

Okay. So the first thing you have to do is expand the right side. (x+4)^2 means (x+4)(x+4). Are you familiar with the distributive property?

OpenStudy (adamk):

I think they mean 2 to the "x+4" power.

OpenStudy (anonymous):

Yes. It isn't x+4 squared though. The equation is \[8=2^{x+4}\]

OpenStudy (adamk):

Well, not exactly sure what you're supposed to do with it. But, I can tell you that there is a transformation of 4 to the right.

OpenStudy (acxbox22):

okay so we know 2^3=8 right?

OpenStudy (anonymous):

How did you get 2^3=8 when the equation is 8=2^(x+4)?

OpenStudy (anonymous):

Wait.. I read that wrong. Sorry

OpenStudy (anonymous):

Yes

OpenStudy (acxbox22):

i am getting there....

OpenStudy (acxbox22):